1/ Find weather the two function are invertible or not, if it is find out its inverse (f^1(x))

1. f:[-π/2, π/2]→[-1,1]; f(x)=sin x
2. f:[0,π]→[-2,2]; f(x)=2cos x

2/ The function f(x)=5/9(x-32) converts Fahrenheit temperatures into Celsius, what is the opposite function for the opposite conversion?
1
Expert's answer
2020-05-27T16:22:40-0400

1.1) Given function "f(x)=sin(x)" is one-one in the given domain and "f(-\pi/2) = -1, f(\pi/2) = 1" so given function is onto also.

Hence, given function is invertible.

Now, "f(x) = sin(x) \implies y = sin(x) \implies x = sin^{-1} (y)".

So, Inverse function is "g(x) = sin^{-1}(x)" , since "(fog)(x) = (gof)(x) = x" .


1.2) Given function is one-one in the given domain and "f(0) = 2, f(\pi) = -2", hence the given function is onto also. Hence given function is onto.

Now, "y = 2cos(x) \implies x = cos^{-1}(y/2)".

So, Inverse function is "g(x) = cos^{-1}(x\/2)" since "(fog)(x) = (gof)(x) = x ."


2) Given "f(x) = (5/9)(x-32) \implies y = (5/9)(x-32)"

"\implies x = (\frac{9}{5}y)+32".

Hence, Opposite function for the opposite conversion is "g(x) = \frac{9}{5} x +32."